# Random currents for signed interactions in the one-dimensional heterogeneous Ising model

### v1

@@ 2.2. Statement of main results
(2.5) $\begin{split}\tilde{\mathbf{T}}_{\{i\cdots j\}}:=\begin{bmatrix}e^{h_{j}}&0\\ 0&e^{-h_{j}}\end{bmatrix}\prod_{x=i}^{j-1}\frac{1}{\sqrt{2\sinh(2|J_{x,x+1}|)}}\begin{bmatrix}e^{J_{x,x+1}}&e^{-J_{x,x+1}}\\ e^{-J_{x,x+1}}&e^{J_{x,x+1}}\end{bmatrix}\begin{bmatrix}e^{h_{x}}&0\\ 0&e^{-h_{x}}\end{bmatrix}\\ \cdot\frac{1}{\sqrt{2\sinh(2|J_{(i,i+1)}|)}}\begin{bmatrix}e^{J_{(i-i+1)}}&e^{-J_{(i,i+1)}}\\ e^{-J_{(i,i+1)}}&e^{J_{(i,i+1)}}\end{bmatrix}\begin{bmatrix}e^{h_{i}}&0\\ 0&e^{-h_{i}}\end{bmatrix},\end{split}$
@@
In the limit where $\operatorname{dist}(i,j)$ is large and the coefficients $\mathbf{J}$ and $\mathbf{h}$ are chosen independently and each identically distributed, it becomes insightful to study the product of matrices that appears in (2.4) as a dynamical system [DH81, DH83, CNPV90, PS92, PPS95, GG22]. For convenience, in this case we can consider $\nu$ to be a restriction of a measure that chooses i.i.d. coefficients on all points and nearest-neighbor edges of $\mathbb{Z}$, which we can also denote $\nu$ by abuse of notation. Because the exact values of quenched covariances will depend slightly on where we restrict the lattice, we will use the notation $\operatorname{cov}_{\{-N,\ldots,N\}}(\cdot)$ where necessary to specify the lattice size.
@@
The upper Lyapunov exponent in Corollary 2.5 can be related to the upper Lyapunov exponent of the matrix in (1.10) via Remark 2.2 and elementary calculation. It may not be possible to express the value of this exponent in closed form. However, its behavior has been studied in detail and is possible to compute numerically to arbitrary precision with efficient algorithms [CNPV90, CPV12, DH81, DH83, PPS95, GG22].
@@ 5. Corollaries of the matrix identity
(5.2) $\left\|M\right\|$ $\leq\prod_{x=0}^{N}\left\|\begin{bmatrix}\sqrt{\coth|J_{x,x+1}|}&0\\ 0&\operatorname{sign}(J_{x,x+1})\sqrt{\tanh|J_{x,x+1}|}\end{bmatrix}\right\|\left\|\begin{bmatrix}\cosh(h_{x})&\sinh(h_{x})\\ \sinh(h_{x})&\cosh(h_{x})\end{bmatrix}\right\|$
(5.3) $=\prod_{x=0}^{N}\sqrt{\coth|J_{x,x+1}|}e^{|h_{x}|}.$
@@ Acknowledgements
JK acknowledges the financial support of the German Academic Exchange Service (DAAD) as well as the Leibniz Association within the Leibniz Junior Research Group on Probabilistic Methods for Dynamic Communication Networks as part of the Leibniz Competition.
@@ Appendix A Calculations for the truncated chain
(B.9) $Z^{\prime}_{++}Z^{\prime}_{--}-Z^{\prime}_{+-}Z^{\prime}_{-+}=\prod_{x=i}^{j-1}2\sinh(2J^{\prime}_{x,x+1}).$
@@
(B.12) $\tilde{\mathbf{T}^{\prime}}_{\{i\cdots j\}}:=\begin{bmatrix}e^{h^{\prime}_{j}}&0\\ 0&e^{-h^{\prime}_{j}}\end{bmatrix}\prod_{x=i}^{j-1}\frac{1}{\sqrt{2\sinh(2|J^{\prime}_{x,x+1}|)}}\begin{bmatrix}e^{J^{\prime}_{x,x+1}}&e^{-J^{\prime}_{x,x+1}}\\ e^{-J^{\prime}_{x,x+1}}&e^{J^{\prime}_{x,x+1}}\end{bmatrix}\begin{bmatrix}e^{h^{\prime}_{x}}&0\\ 0&e^{-h^{\prime}_{x}}\end{bmatrix}.$

### v2

@@ 2.2. Statement of main results
(2.5) $\begin{split}\tilde{\mathbf{T}}_{\{i\cdots j\}}:=\begin{bmatrix}e^{h_{j}}&0\\ 0&e^{-h_{j}}\end{bmatrix}\prod_{x=i}^{j-1}\frac{1}{\sqrt{2\sinh(2|J_{(x,x+1)}|)}}\begin{bmatrix}e^{J_{(x,x+1)}}&e^{-J_{(x,x+1)}}\\ e^{-J_{(x,x+1)}}&e^{J_{(x,x+1)}}\end{bmatrix}\begin{bmatrix}e^{h_{x}}&0\\ 0&e^{-h_{x}}\end{bmatrix}\\ \cdot\frac{1}{\sqrt{2\sinh(2|J_{(i,i+1)}|)}}\begin{bmatrix}e^{J_{(i-i+1)}}&e^{-J_{(i,i+1)}}\\ e^{-J_{(i,i+1)}}&e^{J_{(i,i+1)}}\end{bmatrix}\begin{bmatrix}e^{h_{i}}&0\\ 0&e^{-h_{i}}\end{bmatrix},\end{split}$
@@
In the limit where $\operatorname{dist}(i,j)$ is large and the coefficients $\mathbf{J}$ and $\mathbf{h}$ are chosen independently and each identically distributed, it becomes insightful to study the product of matrices that appears in (2.4) as a dynamical system [DH81, DH83, CNPV90, PS92, PPS95, GG22, CGGH25]. For convenience, in this case we can consider $\nu$ to be a restriction of a measure that chooses i.i.d. coefficients on all points and nearest-neighbor edges of $\mathbb{Z}$, which we can also denote $\nu$ by abuse of notation. Because the exact values of quenched covariances will depend slightly on where we restrict the lattice, we will use the notation $\operatorname{cov}_{\{-N,\ldots,N\}}(\cdot)$ where necessary to specify the lattice size.
@@
The upper Lyapunov exponent in Corollary 2.5 can be related to the upper Lyapunov exponent of the matrix in (1.10) via Remark 2.2 and elementary calculation. It may not be possible to express the value of this exponent in closed form. However, its behavior has been studied in detail and is possible to compute numerically to arbitrary precision with efficient algorithms [CNPV90, CPV12, DH81, DH83, PPS95, GG22, CGGH25].
@@ 5. Corollaries of the matrix identity
(5.2) $\left\|M\right\|$ $\leq\prod_{x=0}^{N}\left\|\begin{bmatrix}\sqrt{\coth|J_{(x,x+1)}|}&0\\ 0&\operatorname{sign}(J_{(x,x+1)})\sqrt{\tanh|J_{(x,x+1)}|}\end{bmatrix}\right\|\left\|\begin{bmatrix}\cosh(h_{x})&\sinh(h_{x})\\ \sinh(h_{x})&\cosh(h_{x})\end{bmatrix}\right\|$
(5.3) $=\prod_{x=0}^{N}\sqrt{\coth|J_{(x,x+1)}|}e^{|h_{x}|}.$
@@ Acknowledgements
JK gratefully acknowledges the financial support of the German Academic Exchange Service (DAAD) as well as the Leibniz Association within the Leibniz Junior Research Group on Probabilistic Methods for Dynamic Communication Networks as part of the Leibniz Competition.
MG has been supported by NSF awards DMS-1954343 and DMS-2348113.
@@ Appendix A Calculations for the truncated chain
(B.9) $Z^{\prime}_{++}Z^{\prime}_{--}-Z^{\prime}_{+-}Z^{\prime}_{-+}=\prod_{x=i}^{j-1}2\sinh(2J^{\prime}_{(x,x+1)}).$
@@
(B.12) $\tilde{\mathbf{T}^{\prime}}_{\{i\cdots j\}}:=\begin{bmatrix}e^{h^{\prime}_{j}}&0\\ 0&e^{-h^{\prime}_{j}}\end{bmatrix}\prod_{x=i}^{j-1}\frac{1}{\sqrt{2\sinh(2|J^{\prime}_{(x,x+1)}|)}}\begin{bmatrix}e^{J^{\prime}_{(x,x+1)}}&e^{-J^{\prime}_{(x,x+1)}}\\ e^{-J^{\prime}_{(x,x+1)}}&e^{J^{\prime}_{(x,x+1)}}\end{bmatrix}\begin{bmatrix}e^{h^{\prime}_{x}}&0\\ 0&e^{-h^{\prime}_{x}}\end{bmatrix}.$